From the given table calculate:

  1. Tax on capital gain at 20%

Firm A

Firm B

Next year’s price

$140.50

$140.50

Dividend

$ 12

$ 13.00

Total pretax payoff

$140.50

$140.50

Today’s stock price

$300

$ 95.78

Capital gain

$ 13.50

$ 7.78
Answer #1

Firm A:

Capital Gain given is = $13.5

Tax @ 20% = 13.5 * 20 / 100 = $2.7

Firm B:

Capital Gain given is = $7.78

Tax @ 20% = 7.78 * 20 / 100 = $1.556